### v2

# The time-dependent quantum harmonic oscillator: a pedagogical approach via the Lewis-Riesenfeld dynamical invariant method

Abstract
In quantum mechanics courses, students often solve the Schrödinger equation for the harmonic oscillator with time-independent parameters. However, time-dependent quantum harmonic oscillators are relevant in modeling several problems as, for instance, the description of quantum motion of particles in traps, shortcuts to adiabaticity, as well as quantum scalar fields evolving in expanding universes. In the present paper, we discuss, with a pedagogical approach, the quantum harmonic oscillator with time-dependent frequency via the Lewis-Riesenfeld dynamical invariant method, revisiting the main steps to obtain the wave function associated with this model, and briefly discussing the relation between this oscillator and the generation of squeezed states. As examples of didactic applications of time-dependent harmonic oscillators and the Lewis-Riesenfeld method in quantum mechanics courses, we solve the following problems: the calculation of the transition probability associated with a harmonic oscillator which undergoes jumps in its frequency, and the analysis of the dynamics of a quantum particle in a Paul trap.
I Introduction
The harmonic oscillator is one of the fundamental systems in physics, permeating from classical mechanics [1, 2, 3, 4] to the quantum field theory [5, 6]. In the context of undergraduate courses, the harmonic oscillator with time-independent parameters is vastly addressed in books in classical (e.g., in Refs. [1, 2, 3]) and in quantum mechanics (e.g., in Refs. [7, 8, 9]). Regarding the harmonic oscillator with time-dependent parameters, here denominated as time-dependent harmonic oscillator (TDHO), although it is addressed in relatively few books of classical mechanics (e.g., in Ref. [1, 4]), it is still commonly discussed in classrooms [10]. On the other hand, the TDHO is more rarely covered in quantum mechanics textbooks (e.g., Refs. [9, 11] include a problem involving the calculation of the transition probability of an oscillator that experiences a sudden jump in its frequency).
While the energy is a constant of motion for the harmonic oscillator with time-independent parameters, it is not for the TDHO. However, Lewis [12, 13] found a certain constant of motion for the TDHO, which is nowadays known as the Lewis’ dynamical invariant. Lutzky, in Ref. [14], showed how to obtain such Lewis’ invariant from Noether’s theorem. Considering that Lutzky’s work was mostly known by specialists, but not by teachers, and also that the TDHO is an important issue in physics education, Abe et al. [10] reviewed Lutzky’s work in a didactic manner, interconnecting, in a simple way, the TDHO, the concept of invariant, and Noether’s theorem.
In the context of quantum mechanics, invariant operators, which can be viewed as generalizations of constants of the motion [15], are useful for finding exact solutions of the Schrödinger equation for systems whose Hamiltonians explicitly depend on time. As an example, one can cite the method proposed by Lewis and Riesenfeld [12, 13, 16], which uses invariant operators to solve the Schrödinger equation for generic time-dependent Hamiltonians and has already been applied in several problems in physics, including those which can be described by a Hamiltonian of a TDHO [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. We can also cite the method discussed by Andrews in Ref. [15], which obtained invariant operators similar to those used by Lewis and Reisenfeld [16], and, with a didactic approach, used them to solve the Schrödinger equation for time-dependent Hamiltonians that are quadratic in position and momentum, with the case of a Hamiltonian describing a TDHO being a particular application.
It is well known that TDHOs are important in quantum physics, since they are relevant in modeling several problems [29, 16, 30, 31, 32, 33]. For example, TDHOs appear in the description of the interaction between a spinless charged quantum particle and a time-dependent external classical electromagnetic field [34, 35, 36], in quantum motion of a particles in traps [37, 38, 39, 40, 41, 42, 43], in the quantization of the free electromagnetic field in nonstationary media [23, 44, 45, 46, 47], in the analysis of scalar fields that evolve in Friedmann-Robertson-Walker metrics [48], in shortcuts to adiabaticity [49, 24, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60], in the generation of squeezed states [61, 62, 63, 64, 65, 66], among others. It is therefore important to improve discussions on TDHOs in the context of undergraduate quantum mechanics courses, an effort that has already begun in some articles in physics teaching journals, such as those already mentioned [10, 15, 33], and also in Refs. [67, 68].
In the present paper, aiming to contribute to the improvement of the discussions on TDHOs, especially in quantum mechanics courses, we didactically discuss the Lewis-Riesenfeld (LR) method [12, 13, 16] and its application to the case of a harmonic oscillator with time-dependent frequency, describing the step-by-step use of the method. As part of our didactic strategy, we map these steps into those already used by undergraduate students to solve the Schrödinger equation for the time-independent harmonic oscillator. As particular applications, we discuss the calculation of the transition probability associated with a TDHO which undergoes a sudden jump in its frequency (the mentioned problem found in Refs. [9, 11]), and we briefly discuss the dynamics of a quantum particle in a Paul trap, which is also modeled by a TDHO with a time-dependent frequency (from the classical point of view, the problem of particles in a Paul trap has been investigated in several physics teaching articles [69, 70, 71, 72, 73, 74, 75, 76, 77]). These calculations can be used as didactic examples of applications of the TDHOs and the LR method in quantum mechanics courses.
This paper is organized as follows. In Sec. II, to prepare the reader for the subsequent sections, we briefly review the procedures for obtaining the wave function associated with the quantum harmonic oscillator with time-independent parameters, as is usually done in quantum mechanics textbooks. In Sec. III, we provide a detailed pedagogical review of the LR method for a generic Hamiltonian $\hat{H}(t)$. In Sec. IV, we provide a review of the application of the LR method to the case of a TDHO with a generic time-dependent frequency. More specifically: In Sec. IV.1, we review the calculation of the wave function (via the LR method) associated with this model. In Sec. IV.2, we discuss the generation of squeezed states by a TDHO with time-dependent frequency. In Sec. IV.3, we use these concepts to solve the problem of the calculation of the transition probability associated to a quantum harmonic oscillator that undergoes a sudden jump in its frequency. In Sec. IV.4, we also use all these tools to discuss the problem of the dynamics of a quantum particle in a Paul trap. In Sec. V, we present our final remarks.
II A brief review of the time-independent quantum harmonic oscillator
As will be seen later, although the solution of the time-dependent quantum harmonic oscillator needs a more robust method to lead with, it has several similarities with the time-independent correspondent problem. In this way, we start by reviewing the problem of the time-independent quantum harmonic oscillator, which is a system that is widely studied in quantum mechanics courses [9, 7].
First, we consider the Schrödinger equation for the one-dimensional time-independent harmonic oscillator,
$i\hbar\frac{\partial}{\partial t}|\psi^{(0)}(t)\rangle=\hat{H}_{0}|\psi^{(0)}(t)\rangle,$ (1)
with the Hamiltonian operator defined by
$\hat{H}_{0}=\frac{\hat{p}^{2}}{2m_{0}}+\frac{1}{2}m_{0}\omega_{0}^{2}\hat{x}^{2},$ (2)
where $m_{0}$ and $\omega_{0}$ are the time-independent mass and angular frequency, respectively. The position and momentum operators, given by $\hat{x}$ and $\hat{p}=-i\hbar\partial/\partial x$ (in position space), satisfy the canonical commutation relation $[\hat{x},\hat{p}]=i\hbar$ [9, 7]. Besides this, it is straight to note that, for the case of a time-independent harmonic oscillator, we have that $\hat{H}_{0}$ is an invariant operator, i.e., $d\hat{H}_{0}/dt=0$.
The solution of Eq. (1) in the position space, given by $\Psi^{(0)}(x,t)=\langle x|\psi^{(0)}(t)\rangle$, can be written as [9, 7]
$\Psi^{(0)}(x,t)=\sum_{n=0}^{\infty}C_{n}^{(0)}\Psi_{n}^{(0)}(x,t),$ (3)
where $\Psi_{n}^{(0)}(x,t)$ are the wave functions associated to each allowed value of energy $(E_{n})$ of the system, and the coefficients $C_{n}^{(0)}$ depends only on the initial conditions. Using this solution into Eq. (1) in the position space, and performing the method of separation of variables, one finds [9, 7]
$\Psi_{n}^{(0)}(x,t)=\Phi_{n}^{(0)}(x)\exp\left(-\frac{iE_{n}t}{\hbar}\right),$ (4)
such that $E_{n}$ and $\Phi_{n}^{(0)}(x)$ have to be obtained by means of the eigenvalue equations
$\hat{H}_{0}\Phi_{n}^{(0)}(x)=E_{n}\Phi_{n}^{(0)}(x).$ (5)
For our purposes, here we solve this equation by means of an algebraic method, using the so-called annihilation and creation operators (sometimes they are also denominated as lowering and raising operators, respectively).
The algebraic method for solving Eq. (5) can be performed by introducing the annihilation ($\hat{a}_{0}$) and creation ($\hat{a}_{0}^{\dagger}$) operators, which are defined in terms of the operators $\hat{x}$ and $\hat{p}$ as [9, 7]
$\hat{a}_{0}=$ $\;\sqrt{\frac{m_{0}\omega_{0}}{2\hbar}}\left(\hat{x}+\frac{i\hat{p}}{m_{0}\omega_{0}}\right),$ (6)
$\hat{a}_{0}^{\dagger}=$ $\;\sqrt{\frac{m_{0}\omega_{0}}{2\hbar}}\left(\hat{x}-\frac{i\hat{p}}{m_{0}\omega_{0}}\right),$ (7)
where one notes that these operators are non-Hermitian. From these expressions and from the commutation relation between $\hat{x}$ and $\hat{p}$, one can obtain that $[\hat{a}_{0},\hat{a}_{0}^{\dagger}]=1$, and we can rewrite the operator $\hat{H}_{0}$ as
$\hat{H}_{0}=\hbar\omega_{0}\left(\hat{a}_{0}^{\dagger}\hat{a}_{0}+\frac{1}{2}\right).$ (8)
The product $\hat{a}_{0}^{\dagger}\hat{a}_{0}$ in this equation is denominated as the number operator $\hat{N}_{0}=\hat{a}_{0}^{\dagger}\hat{a}_{0}$. In this way, let us denote the eigenvalues of $\hat{N}_{0}$ as $n$ and its normalized eigenstates as $\left|n\right\rangle$ [whose representation in the position space, $\left\langle x|n\right\rangle$, is the function $\Phi_{n}^{(0)}(x)$], so that
$\hat{N}_{0}|n\rangle=n|n\rangle.$ (9)
From this, one can obtain that
$\hat{H}_{0}|n\rangle=\hbar\omega_{0}\left(n+\frac{1}{2}\right)|n\rangle,$ (10)
from which we can conclude that the energy eigenvalues (that is, the eigenvalues of the operator $\hat{H}_{0}$) are given by [9, 7]
$E_{n}=\hbar\omega_{0}\left(n+\frac{1}{2}\right).$ (11)
To calculate the eigenstates $\left|n\right\rangle$, one has to obtain the action of the operators $\hat{a}_{0}$ and $\hat{a}_{0}^{\dagger}$ in this states, which, from the commutation relation between these operators and Eq. (9), one finds that $\hat{a}_{0}$ and $\hat{a}_{0}^{\dagger}$ satisfy the following algebra [9, 7]:
$\hat{a}_{0}|n\rangle=$ $\;\sqrt{n}|n-1\rangle,$ (12)
$\hat{a}_{0}^{\dagger}|n\rangle=$ $\;\sqrt{n+1}|n+1\rangle,$ (13)
which justify the nomenclature annihilation and creation operators (or lowering and raising operators, respectively). Note that repeatedly application of $\hat{a}_{0}$ produces states with smaller and smaller values of $n$. In this way, the action of the operator $\hat{a}_{0}$ in the fundamental eigenstate $|0\rangle$ has to be null, i.e.,
$\hat{a}_{0}|0\rangle=0.$ (14)
From the fundamental eigenstate, we can generate any excited state $|n\rangle$ by applying $\hat{a}_{0}^{\dagger}$ in it repeatedly. Thus, let us calculate it by applying $\langle x|$ in Eq. (14), and using Eq. (6) in position space, so that one obtains
$\left(x+\frac{\hbar}{m_{0}\omega_{0}}\frac{\partial}{\partial x}\right)\Phi_{0}^{(0)}(x)=0,$ (15)
where $\Phi_{0}^{(0)}(x)=\langle x|0\rangle$. By solving this differential equation, one obtains the normalized solution [9, 7]
$\Phi_{0}^{(0)}(x)=\left(\frac{m_{0}\omega_{0}}{\pi\hbar}\right)^{\frac{1}{4}}\exp\left(-\frac{m_{0}\omega_{0}x^{2}}{2\hbar}\right).$ (16)
By applying $\hat{a}_{0}^{\dagger}$ in $\Phi_{0}^{(0)}(x)$ repeatedly, one obtains the following expression for $\Phi_{n}^{(0)}(x)$:
$\Phi_{n}^{(0)}(x)=\frac{(a_{0}^{\dagger})^{n}}{\sqrt{n!}}\Phi_{0}^{(0)}(x),$ (17)
which, using Eqs. (7) (in position space) and (16), one obtains that [9, 7]
$\Phi_{n}^{(0)}(x)=\frac{1}{\sqrt{2^{n}n!}}\Phi_{0}^{(0)}(x){\cal H}_{n}\left(\sqrt{\frac{m_{0}\omega_{0}}{\hbar}}x\right),$ (18)
where ${\cal H}_{n}$ are the usual Hermite polynomials of order $n$. In this sense, using Eqs. (11) and (18) in Eq. (4), we obtain the wave function of the time-independent quantum harmonic oscillator [7, 9]:
$\Psi_{n}^{(0)}(x,t)=\frac{1}{\sqrt{2^{n}n!}}\left(\frac{m_{0}\omega_{0}}{\pi\hbar}\right)^{\frac{1}{4}}\exp\left[-i\left(n+\frac{1}{2}\right)\omega_{0}t\right]\exp\left(-\frac{m_{0}\omega_{0}x^{2}}{2\hbar}\right){\cal H}_{n}\left(\sqrt{\frac{m_{0}\omega_{0}}{\hbar}}x\right).$ (19)
In summary, the algebraic method used to find the solution of Eq. (5) consists in the following steps:
  1. 1.
    Define the annihilation $(\hat{a}_{0})$ and creation $(\hat{a}_{0}^{\dagger})$ operators [Eqs. (6) and (7), respectively];
  2. 2.
    Express the Hamiltonian operator $\hat{H}_{0}$ in terms of $\hat{a}_{0}$ and $\hat{a}_{0}^{\dagger}$ [Eq. (8)];
  3. 3.
    Define the number operator $\hat{N}=\hat{a}_{0}^{\dagger}\hat{a}_{0}$ with eigenvalue and eigenstate denoted by $n$ and $\left|n\right\rangle$, respectively [Eq. (9)];
  4. 4.
    Find the effect of $\hat{a}_{0}$ and $\hat{a}_{0}^{\dagger}$ on the state $\left|n\right\rangle$ [Eqs. (12) and (13)];
  5. 5.
    Define the ground state $\left|0\right\rangle$ from the property that the application of the annihilation operator in this state has to be null [Eq. (14)], and calculate $\Phi_{0}^{(0)}(x)=\langle x|0\rangle$;
  6. 6.
    Calculate the excited states $\Phi_{n}^{(0)}(x)$ by applying repeatedly the creation operator in $\Phi_{0}^{(0)}(x)$ [Eq. (17)];
  7. 7.
    Find the general solution $\Psi^{(0)}(x,t)$ of the Schrödinger equation [Eq. (3), with $\Psi_{n}^{(0)}(x,t)$ given in Eq. (19)].
We highlight that, for a TDHO [where $\hat{H}_{0}\to\hat{H}(t)$], the LR method prescribes that the wave function can be found by means of a sequence of steps similar to these ones used in the case of a time-independent harmonic oscillator. In the time-independent case, $\hat{H}_{0}$ is an invariant operator, since $d\hat{H}_{0}/dt=0$. In the time-dependent case, the Hamiltonian is not an invariant operator anymore, but, according to the LR method, we can find another operator, denoted by $\hat{I}(t)$, so that this new operator is invariant, i.e., $d\hat{I}(t)/dt=0$. Thus, by finding $\hat{I}(t)$, the next steps are similar to those shown in Eqs. (3)-(19), as discussed next.
III Lewis-Riesenfeld dynamical invariant method in quantum mechanics
Before discussing the solution of a TDHO, let us start reviewing the well known LR dynamical invariant method [12, 13, 16], which consists of a mechanism for finding exact solutions of the Schrödinger equation for systems whose Hamiltonians explicitly depend on time, i.e., for finding solutions of the following equation:
$i\hbar\frac{\partial}{\partial t}|\psi(t)\rangle=\hat{H}(t)|\psi(t)\rangle.$ (20)
The starting point for solving this equation consists in finding a Hermitian operator $\hat{I}(t)$, called invariant operator. Such operator has the property of being a constant of motion, i.e., $d\hat{I}(t)/dt=0$, such that, from the Heisenberg equation [16, 7], one has
$\frac{\partial\hat{I}(t)}{\partial t}+\frac{1}{i\hbar}\bigl{[}\hat{I}(t),\hat{H}(t)\bigr{]}=0.$ (21)
According to Ref. [16]: “The action of the invariant operator on a Schrödinger state vector produces another solution of the Schrödinger equation. This result is valid for any invariant, even if the latter involves the operation of time differentiation”. To see this, one has to apply the operator $\partial\hat{I}(t)/\partial t$ on the state $|\psi(t)\rangle$, so that, by using Eq. (21), one obtains
$\frac{\partial\hat{I}(t)}{\partial t}|\psi(t)\rangle=-\frac{1}{i\hbar}\bigl{[}\hat{I}(t)\hat{H}(t)-\hat{H}(t)\hat{I}(t)\bigr{]}|\psi(t)\rangle.$ (22)
Using Eq. (20) in this equation, one sees that
$i\hbar\frac{\partial}{\partial t}\bigl{[}\hat{I}(t)|\psi(t)\rangle\bigr{]}=\hat{H}(t)\bigl{[}\hat{I}(t)|\psi(t)\rangle\bigr{]},$ (23)
which shows that $\hat{I}(t)|\psi(t)\rangle$ is also a solution of the Schrödinger equation.
Following Ref. [16]: “We assume that the invariant operator is one of a complete set of commuting observables, so that there is a complete set of eigenstates of $\hat{I}(t)$”. In this way, let us denote the eigenvalues of $\hat{I}(t)$ as $\lambda$, which are time-independent (see Appendix A for more details) [16], and its orthonormal eigenstates associated to $\lambda$ as $|\lambda,\zeta;t\rangle$, with $\zeta$ being all quantum numbers, other than $\lambda$, needed to specify the eigenstates. Therefore, one writes (for a discrete spectrum)
$\hat{I}(t)|\lambda,\zeta;t\rangle=$ $\;\lambda|\lambda,\zeta;t\rangle,$ (24)
$\sum_{\lambda,\zeta}|\lambda,\zeta;t\rangle\langle\lambda,\zeta;t|=$  $\;1,$ (25)
$\langle\lambda^{\prime},\zeta^{\prime};t|\lambda,\zeta;t\rangle=$ $\;\delta_{\lambda^{\prime},\lambda}\delta_{\zeta^{\prime},\zeta},$ (26)
where $\delta_{i,j}$ is the Kronecker delta (for a continuous spectrum, this is replaced by Dirac delta functions).
We remark that the eigenstates $|\lambda,\zeta;t\rangle$ do not satisfy, in general, the Schrödinger equation, i.e.,
$i\hbar\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle\neq\hat{H}(t)|\lambda,\zeta;t\rangle.$ (27)
Nevertheless, each state $|\lambda,\zeta;t\rangle$ can be multiplied by a phase factor, resulting in a state $|\lambda,\zeta;t\rangle_{\alpha}$, such that
$|\lambda,\zeta;t\rangle_{\alpha}=\exp[i\alpha_{\lambda,\zeta}(t)]|\lambda,\zeta;t\rangle,$ (28)
with $\alpha_{\lambda,\zeta}(t)$ being a real and time-dependent phase function. Furthermore, following Ref. [16]: “If the invariant does not involve time differentiation, then we are able to derive a simple and explicit rule for choosing the phases of the eigenstates of $\hat{I}(t)$ such that these states themselves satisfy the Schrödinger equation”. In this way, assuming that the invariant operator does not involve time differentiation [16], it is straightforward to see that the states $|\lambda,\zeta;t\rangle_{\alpha}$ are also eigenstates of $\hat{I}(t)$, and form an orthonormal basis. Besides this, if we require that the state vectors $|\lambda,\zeta;t\rangle_{\alpha}$ are solutions of the Schrödinger equation,
$i\hbar\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle_{\alpha}=\hat{H}(t)|\lambda,\zeta;t\rangle_{\alpha},$ (29)
then, one obtains that the phase functions $\alpha_{\lambda,\zeta}(t)$ have to obey the following equation:
$\frac{d\alpha_{\lambda,\zeta}(t)}{dt}=\langle\lambda,\zeta;t|\biggl{[}i\frac{\partial}{\partial t}-\frac{1}{\hbar}\hat{H}(t)\biggr{]}|\lambda,\zeta;t\rangle.$ (30)
The solution $\alpha_{\lambda,\zeta}(t)$ of the equation above is known as the Lewis-Riesenfeld phase and, in turn, is composed of the dynamical and geometric phases. These phases have applications in different areas of physics (see, for example, Refs. [78, 79, 80, 81]).
The original problem of finding the eigenstates and eigenvalues of $\hat{H}(t)$ is replaced by the problem of finding the eigenstates and eigenvalues of $\hat{I}(t)$ and the phase functions $\alpha_{\lambda,\zeta}(t)$. By knowing the operator $\hat{I}(t)$ (which is, in general, obtained from an ansatz), one can calculate its eigenvalues and eigenstates by means of Eq. (24), and, with this, obtain the phase functions $\alpha_{\lambda,\zeta}(t)$ from Eq. (30). Therefore, the general solution $|\psi(t)\rangle$ of Eq. (20) can be described as a linear combination of $|\lambda,\zeta;t\rangle$ as
$|\psi(t)\rangle=\sum_{\lambda,\zeta}C_{\lambda,\zeta}\exp[i\alpha_{\lambda,\zeta}(t)]|\lambda,\zeta;t\rangle,$ (31)
in which the time-independent coefficients $C_{\lambda,\zeta}$ depend on the initial conditions. For more details on the steps between Eqs. (28)-(31), see Appendix B.
In summary, the LR method consists of the following steps:
  1. 1.
    One proposes a Hermitian invariant operator $\hat{I}(t)$, which satisfies Eq. (21), and does not contain time derivatives;
  2. 2.
    From Eq. (24), one calculates the eigenvalues $\lambda$ and eigenstates $|\lambda,\zeta;t\rangle$ associated to $\hat{I}(t)$;
  3. 3.
    Using Eq. (30), one obtains the phase functions $\alpha_{\lambda,\zeta}(t)$;
  4. 4.
    Finally, by means of Eq. (31), one calculates the state vector $|\psi(t)\rangle$.
As a consistency check of Eq. (31), let us discuss the situation in which we have a time-independent Hamiltonian $[\hat{H}(t)=\hat{H}_{\text{TI}}]$. In this situation, one can propose that the invariant operator is given by $\hat{I}(t)=\hat{I}_{\text{TI}}$, with $\hat{I}_{\text{TI}}$ being a time-independent operator. This results in $\partial\hat{I}(t)/\partial t=\partial\hat{I}_{\text{TI}}/\partial t=0$, which, from Eq. (21), leads to $\bigl{[}\hat{I}(t),\hat{H}(t)\bigr{]}=\bigl{[}\hat{I}_{\text{TI}},\hat{H}_{\text{TI}}\bigr{]}=0$. As a consequence, the eigenstates of $\hat{H}_{\text{TI}}$, denoted by $|E_{j}\rangle$, are also eigenstates of $\hat{I}_{\text{TI}}$ [7], which means that the state $|\lambda,\zeta;t\rangle$ coincides with $|E_{j}\rangle$, so that we can do the mapping $|\lambda,\zeta;t\rangle\to|E_{j}\rangle$ in Eq. (31) and write
$|\psi(t)\rangle=\sum_{j}C_{j}\exp[i\alpha_{j}(t)]|E_{j}\rangle.$ (32)
The same mapping has to be made in Eq. (30) in order to obtain the phase function for this case. Since the state $|E_{j}\rangle$ is time-independent $[\partial|E_{j}\rangle/\partial t=0]$ and using the eigenvalue equation $\hat{H}_{\text{TI}}|E_{j}\rangle=E_{j}|E_{j}\rangle$, where $E_{j}$ are the eigenvalues of $\hat{H}_{\text{TI}}$, one can find the solution of Eq. (30), which is given by
$\alpha_{j}(t)=-\frac{E_{j}t}{\hbar}.$ (33)
By substituting this result into Eq. (32), one sees that we recover the general result for $|\psi(t)\rangle$ commonly presented in quantum mechanics textbooks [9, 7] in the context of problems involving time-independent Hamiltonians.
By comparing the characteristics of a time-dependent problem within the context of the LR method with the corresponding time-independent case, we can find correlations, which are summarized in Table 1.
Time-independent case Time-dependent case
Schrödinger equation $i\hbar\partial_{t}|\psi(t)\rangle=\hat{H}_{\text{TI}}|\psi(t)\rangle$ $i\hbar\partial_{t}|\psi(t)\rangle=\hat{H}(t)|\psi(t)\rangle$
Invariant operator $\hat{H}_{\text{TI}}$ $\hat{I}(t)$
Eigenvalue equation $\hat{H}_{\text{TI}}|E_{j}\rangle=E_{j}|E_{j}\rangle$ $\hat{I}(t)|\lambda,\zeta;t\rangle=\lambda|\lambda,\zeta;t\rangle$
General solution $|\psi(t)\rangle=\sum_{j}C_{j}\exp[i\alpha_{j}(t)]|E_{j}\rangle$ $|\psi(t)\rangle=\sum_{\lambda,\zeta}C_{\lambda,\zeta}\exp[i\alpha_{\lambda,\zeta}(t)]|\lambda,\zeta;t\rangle$
[…]
Table 1: Comparison between the time-independent and time-dependent cases.
IV The solution of a quantum harmonic oscillator with time-dependent frequency via Lewis-Riesenfeld dynamical invariant method
In the present section, we analyze the one-dimensional TDHO with time-dependent frequency $\omega(t)$, which is described by the Hamiltonian operator
$\hat{H}(t)=\frac{\hat{p}^{2}}{2m_{0}}+\frac{1}{2}m_{0}\omega(t)^{2}\hat{x}^{2}.$ (34)
For didactic reasons, in this paper we focus only on the case of a time-dependent frequency, but, in the literature, one finds more general models that have been studied. For example, in Refs. [30, 31], one finds the discussion of the case where the frequency and the mass depend on time. In Refs. [82, 83, 84], one finds the discussion of the driven generalized TDHO, where there are time-dependent parameters coupled to linear terms in $\hat{x}$ and $\hat{p}$, as well as external forces, in the Hamiltonian operator. In Ref. [85], one finds the discussion of the case where there are dependences of the type $1/\hat{x}$ and $1/\hat{x}^{2}$, multiplied by time-dependent factors, in the Hamiltonian operator. Finally, in Refs. [86, 87, 88], there are cases where there are damped harmonic oscillators in a time-dependent noncommutative space, and, in Ref. [89], there is the case of systems with two coupled TDHOs.
IV.1 Calculation of the wave function
IV.1.1 The invariant operator
Originally, the invariant operator $\hat{I}(t)$ for a TDHO was obtained from the ansatz [16]
$\hat{I}(t)=\frac{1}{2}[\eta_{1}(t)\hat{x}^{2}+\eta_{2}(t)\hat{p}^{2}+\eta_{3}(t)\{\hat{x},\hat{p}\}],$ (35)
where $\eta_{1}(t)$, $\eta_{2}(t)$ and $\eta_{3}(t)$ are real time-dependent functions [which ensures that $\hat{I}(t)$ is Hermitian], and $\left\{\hat{x},\hat{p}\right\}$ is the anticommutator between the operators $\hat{x}$ and $\hat{p}$. Since the invariant operator must be a constant of motion $[d\hat{I}(t)/dt=0]$, then we require that this ansatz satisfies Eq. (21). Thus, substituting Eq. (35) into Eq. (21), one finds a system of coupled differential equations involving the functions $\eta_{1}(t)$, $\eta_{2}(t)$ and $\eta_{3}(t)$:
$\dot{\eta}_{1}(t)=$  $\;2m_{0}\omega(t)^{2}\eta_{3}(t),$ (36)
$\dot{\eta}_{2}(t)=$ $-\frac{2\eta_{3}(t)}{m_{0}},$ (37)
$\dot{\eta}_{3}(t)=$ $\;m_{0}\omega(t)^{2}\eta_{2}(t)-\frac{\eta_{1}(t)}{m_{0}}.$ (38)
To decouple this system, we introduce a real parameter $\rho(t)$, such that
$\eta_{2}(t)=\rho(t)^{2}.$ (39)
Then, from Eq. (37), we have
$\eta_{3}(t)=-m_{0}\rho(t)\dot{\rho}(t),$ (40)
as well as, from Eq. (38), we obtain
$\eta_{1}(t)=m_{0}^{2}\dot{\rho}(t)^{2}+m_{0}^{2}\rho(t)[\ddot{\rho}(t)+\omega(t)^{2}\rho(t)].$ (41)
Using these expressions for $\eta_{1}(t)$ and $\eta_{3}(t)$ in Eq. (36), one obtains an equation that determines the dynamics of the parameter $\rho(t)$, which is given by
$\ddot{\rho}(t)+\omega(t)^{2}\rho(t)=\frac{1}{m_{0}^{2}\rho(t)^{3}},$ (42)
and is called Ermakov-Pinney equation [90, 91, 92, 93]. Furthermore, from Eqs. (39)-(41), one sees that the functions $\eta_{1}(t)$, $\eta_{2}(t)$ and $\eta_{3}(t)$ are completely determined by the parameter $\rho(t)$. Thus, substituting these equations into Eq. (35), one finds that the invariant operator, associated to a TDHO with time-dependent frequency, is given by [12, 13, 16, 30, 31]
$\hat{I}(t)=\frac{1}{2}\left\{\left[\frac{\hat{x}}{\rho(t)}\right]^{2}+\left[\rho(t)\hat{p}-m_{0}\dot{\rho}(t)\hat{x}\right]^{2}\right\}.$ (43)
In this way, by knowing how the frequency of the oscillator depend on time, one can obtain an expression for the invariant operator by only calculating the parameter $\rho(t)$ using Eq. (42). We highlight that although Eq. (42) is a nonlinear differential equation, it has a general exact solution, which is given by [91, 92, 90]
$\rho(t)=\sqrt{Au(t)^{2}+Bv(t)^{2}+2Cu(t)v(t)},$ (44)
where $u(t)$ and $v(t)$ are two linearly independent solutions of the homogeneous equation
$\ddot{k}(t)+\omega(t)^{2}k(t)=0,$ (45)
and $A$, $B$ and $C$ are integration constants that depend on the boundary conditions of the problem at a certain time instant $t=t_{0}$ and satisfy the relation
$AB-C^{2}=\frac{1}{m_{0}^{2}\,W\left[u\left(t=t_{0}\right),v\left(t=t_{0}\right)\right]^{2}},$ (46)
with $W\left[u\left(t=t_{0}\right),v\left(t=t_{0}\right)\right]$ being the Wronskian of $u(t=t_{0})$ and $v(t=t_{0})$.
IV.1.2 The eigenvalues and eigenfunctions of the invariant operator
Following Ref. [16]: “The eigenstates and eigenvalues of the invariant operator $\hat{I}(t)$ may be found by an operator technique that is completely analogous to the method introduced by Dirac for diagonalizing the Hamiltonian of a constant-frequency harmonic oscillator”. In other words, from this point, the steps for finding the eigenstates and eigenvalues of $\hat{I}(t)$ are similar to those discussed in Sec. II and used to find the eigenstates and eigenvalues of $\hat{H}_{0}$.
We start by defining the operators $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$ as [16]
$\hat{a}(t)=\frac{1}{\sqrt{2\hbar}}\left\{\frac{\hat{x}}{\rho(t)}+i\left[\rho(t)\hat{p}-m_{0}\dot{\rho}(t)\hat{x}\right]\right\},$ (47)
and
$\hat{a}^{\dagger}(t)=\frac{1}{\sqrt{2\hbar}}\left\{\frac{\hat{x}}{\rho(t)}-i\left[\rho(t)\hat{p}-m_{0}\dot{\rho}(t)\hat{x}\right]\right\}.$ (48)
From these expressions, one can obtain that $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$ satisfy the commutation relation $\left[\hat{a}(t),\hat{a}^{\dagger}(t)\right]=1$. By knowing this, the next step is to express the operator $\hat{I}(t)$ [Eq. (43)] in terms of $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$, which one obtains that
$\hat{I}(t)=\hbar\left[\hat{a}^{\dagger}(t)\hat{a}(t)+\frac{1}{2}\right].$ (49)
In the sequence, we define the number operator $\hat{N}(t)=\hat{a}^{\dagger}(t)\hat{a}(t)$, whose eigenvalue and eigenstate are given by $n$ and $|n;t\rangle$, respectively, so that one writes [16]
$\hat{N}(t)|n;t\rangle=n|n;t\rangle,$ (50)
with $n\in\mathbb{N}$. Thus,
$\hat{I}(t)|n;t\rangle=\hbar\left(n+\frac{1}{2}\right)|n;t\rangle.$ (51)
By comparing Eqs. (24) and (51), one sees that the eigenstate of the operator $\hat{I}(t)$ can be simply written as $|n;t\rangle$ [i.e., we can do the mapping $|\lambda,\zeta;t\rangle\to|n;t\rangle$], and its eigenvalues are given by
$\lambda_{n}=\hbar\left(n+\frac{1}{2}\right).$ (52)
The next step is to find the effect of $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$ in the state $|n;t\rangle$, which, from the commutation relation between these operators, one obtains
$\hat{a}(t)|n;t\rangle=$ $\;\sqrt{n}|n-1;t\rangle,$ (53)
$\hat{a}^{\dagger}(t)|n;t\rangle=$ $\;\sqrt{n+1}|n+1;t\rangle.$ (54)
In the sequence, by knowing Eq. (53), we can define the ground state of $\hat{I}(t)$, denoted by $|0;t\rangle$, which has to satisfy
$\hat{a}(t)|0;t\rangle=0.$ (55)
By applying $\langle x|$ in this equation, and using Eq. (47), one can obtain that [13]
$\biggl{\{}\frac{x}{\rho(t)}-i\Bigl{[}i\hbar\rho(t)\frac{\partial}{\partial x}+m_{0}\dot{\rho}(t)x\Bigr{]}\biggr{\}}\Phi_{0}(x,t)=0,$ (56)
where $\Phi_{0}(x,t)=\langle x|0;t\rangle$, whose solution is given by the Gaussian distribution
$\Phi_{0}(x,t)=\biggl{[}\frac{1}{\pi\hbar\rho(t)^{2}}\biggr{]}^{\frac{1}{4}}\exp\left\{\frac{im_{0}}{2\hbar}\left[\frac{\dot{\rho}(t)}{\rho(t)}+\frac{i}{m_{0}\rho(t)^{2}}\right]x^{2}\right\}.$ (57)
From this solution, we can calculate the excited states $\Phi_{n}(x,t)=\langle x|n;t\rangle$ by applying repeatedly the operator $\hat{a}^{\dagger}(t)$ in $\Phi_{0}(x,t)$, so that one finds
$\Phi_{n}(x,t)=\frac{\bigl{[}a(t)^{\dagger}\bigr{]}^{n}}{\sqrt{n!}}\Phi_{0}(x,t).$ (58)
Using Eqs. (48) and (57) into Eq. (58), one obtains the eigenfunctions of $\hat{I}(t)$ (see Appendix C for more details) [30, 31]:
$\Phi_{n}(x,t)=\frac{1}{\sqrt{2^{n}n!}}\Phi_{0}(x,t){\cal H}_{n}\left[\frac{x}{\hbar^{\frac{1}{2}}\rho(t)}\right].$ (59)
Note that, like the operator $\hat{I}(t)$, the eigenfunctions $\Phi_{n}(x,t)$ are completely determined if we know the parameter $\rho(t)$. Moreover, the eigenfunctions $\Phi_{n}(x,t)$ clearly form a complete orthonormal set (as required), since the Hermite polynomials constitute a basis of Hilbert space [94]. The last step is to find the general solution of the Schrödinger equation [Eq. (31)]. For this, we have to calculate the phase functions $\alpha_{\lambda,\zeta}(t)$ for this case by means of Eq. (30).
IV.1.3 The phase functions
For the TDHO, we can do the mapping $\alpha_{\lambda,\zeta}(t)\to\alpha_{n}(t)$ in Eq. (30), so that it can be written as
$\frac{d\alpha_{n}(t)}{dt}=\langle n;t|\biggl{[}i\frac{\partial}{\partial t}-\frac{1}{\hbar}\hat{H}(t)\biggr{]}|n;t\rangle.$ (60)
Let us manipulate each term on the right side of the equation separately. Focusing on the first term, $i\langle n;t|\frac{\partial}{\partial t}|n;t\rangle$, by writing $|n;t\rangle=[\hat{a}^{\dagger}(t)/\sqrt{n}\;]|n-1;t\rangle$ and performing the time derivative, one obtains
$\langle n;t|\frac{\partial}{\partial t}|n;t\rangle=\frac{1}{\sqrt{n}}\langle n;t|\frac{\partial\hat{a}^{\dagger}(t)}{\partial t}|n-1;t\rangle+\langle n-1;t|\frac{\partial}{\partial t}|n-1;t\rangle.$ (61)
Moreover, by performing a time derivative in Eq. (48) and substituting the result into Eq. (61), one obtains
$\langle n;t|\frac{\partial}{\partial t}|n;t\rangle=\frac{i}{2}m_{0}[\rho(t)\ddot{\rho}(t)-\dot{\rho}(t)^{2}]+\langle n-1;t|\frac{\partial}{\partial t}|n-1;t\rangle.$ (62)
Using recursively this equation one obtains, by induction, that [16]
$\langle n;t|\frac{\partial}{\partial t}|n;t\rangle=\frac{in}{2}m_{0}[\rho(t)\ddot{\rho}(t)-\dot{\rho}(t)^{2}]+\langle 0;t|\frac{\partial}{\partial t}|0;t\rangle.$ (63)
Inserting an identity operator in the term $\langle 0;t|\frac{\partial}{\partial t}|0;t\rangle$, and using Eq. (57), one obtains (see Appendix D for more details)
$\langle 0;t|\frac{\partial}{\partial t}|0;t\rangle=\frac{im_{0}}{4}[\rho(t)\ddot{\rho}(t)-\dot{\rho}(t)^{2}],$ (64)
which, substituting in Eq. (63), results in
$i\langle n;t|\frac{\partial}{\partial t}|n;t\rangle=-\frac{1}{2}\left(n+\frac{1}{2}\right)m_{0}\bigl{[}\rho(t)\ddot{\rho}(t)-\dot{\rho}(t)^{2}\bigr{]}.$ (65)
Now, let us focus on the second term on the right side of Eq. (60), $\langle n;t|\hat{H}(t)|n;t\rangle$. From Eq. (34), we have
$\langle n;t|\hat{H}(t)|n;t\rangle=\frac{1}{2m_{0}}\langle n;t|\hat{p}^{2}|n;t\rangle+\frac{1}{2}m_{0}\omega(t)^{2}\langle n;t|\hat{x}^{2}|n;t\rangle.$ (66)
From Eqs. (47) and (48), we can express the operators $\hat{x}$ and $\hat{p}$ in terms of $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$, and obtain that
$\langle n;t|\hat{x}^{2}|n;t\rangle=$ $\left(n+\frac{1}{2}\right)\hbar\rho(t)^{2},$ (67)
$\langle n;t|\hat{p}^{2}|n;t\rangle=$ $\left(n+\frac{1}{2}\right)\hbar\bigl{[}m_{0}^{2}\dot{\rho}(t)^{2}+\rho(t)^{-2}\bigr{]},$ (68)
which implies in
$\langle n;t|\hat{H}(t)|n;t\rangle=\frac{\hbar}{2m_{0}}\left(n+\frac{1}{2}\right)\bigl{[}\rho(t)^{-2}+m_{0}^{2}\dot{\rho}(t)^{2}+m_{0}^{2}\omega(t)^{2}\rho(t)^{2}\bigr{]}.$ (69)
Thus, substituting Eqs. (65) and (69) into Eq. (60), and using Eq. (42), one obtains that the phase function $\alpha_{n}(t)$ is given by [12, 16, 30, 31]:
$\alpha_{n}(t)=-\left(n+\frac{1}{2}\right)\int_{0}^{t}\frac{dt^{\prime}}{m_{0}\rho\left(t^{\prime}\right)^{2}}.$ (70)
IV.1.4 The general solution of the Schrödinger equation
For the TDHO, Eq. (31) is rewritten as
$|\psi(t)\rangle=\sum_{n=0}^{\infty}C_{n}\exp[i\alpha_{n}(t)]|n;t\rangle.$ (71)
By applying $\langle x|$ in both sides of this equation, one obtains
$\Psi(x,t)=\sum_{n=0}^{\infty}C_{n}\Psi_{n}(x,t),$ (72)
where $\Psi(x,t)=\langle x|\psi(t)\rangle$ and
$\Psi_{n}(x,t)=\exp[i\alpha_{n}(t)]\Phi_{n}(x,t).$ (73)
Substituting Eqs. (59) and (70) in this equation, one finally finds the wave function $\Psi_{n}(x,t)$, which is given by [30, 31, 32, 95]
$\Psi_{n}(x,t)=$ $\;\frac{1}{\sqrt{2^{n}n!\pi^{\frac{1}{2}}\hbar^{\frac{1}{2}}\rho(t)}}\exp\left[-i\left(n+\frac{1}{2}\right)\int_{0}^{t}\frac{dt^{\prime}}{m_{0}\rho\left(t^{\prime}\right)^{2}}\right]$
×$\times\exp\left\{\frac{im_{0}}{2\hbar}\left[\frac{\dot{\rho}(t)}{\rho(t)}+\frac{i}{m_{0}\rho(t)^{2}}\right]x^{2}\right\}{\cal H}_{n}\left[\frac{x}{\hbar^{\frac{1}{2}}\rho(t)}\right].$ (74)
One can conclude that the initial problem is reduced to simply obtain the solution of the parameter $\rho(t)$ by means of Eq. (42), which, although is a nonlinear differential equation, has an exact solution. This is equivalent to solving a classical harmonic oscillator with time-dependent frequency, as Husimi concluded in 1953 [29].
IV.1.5 Static case
For the particular case where the mass and the frequency of the oscillator are independent of time [$m(t)=m_{0}$ and $\omega(t)=\omega_{0}$], the parameter $\rho(t)$ takes the form [16, 31, 32]
$\rho(t)=\rho_{0}=\frac{1}{\sqrt{m_{0}\omega_{0}}}.$ (75)
In this case, it is immediate to see that the Hamiltonian $\hat{H}(t)$ falls back to the Hamiltonian $\hat{H}_{0}$ of a time-independent oscillator, and that the invariant operator becomes $\hat{I}(t)=\hat{I}_{0}=\hat{H}_{0}/\omega_{0}$ [16], and the operators $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$ become, respectively, $\hat{a}_{0}$ and $\hat{a}_{0}^{\dagger}$ [Eqs. (6) and (7)]. Thus, the wave function ${\Psi_{n}(x,t)}$ recover the wave function ${\Psi_{n}^{\left(0\right)}(x,t)}$ related to the time-independent oscillator, given by Eq. (19), as expected. Similar to the previous section, by comparing the characteristics of a TDHO problem within the context of the LR method with the corresponding time-independent case, we can find correlations, which are summarized in Table 2.
Time-independent case Time-dependent case
Schrödinger equation $i\hbar\partial_{t}|\psi(t)\rangle=\hat{H}_{0}|\psi(t)\rangle$ $i\hbar\partial_{t}|\psi(t)\rangle=\hat{H}(t)|\psi(t)\rangle$
Invariant operator $\hat{H}_{0}$ given in Eq. (2) $\hat{I}(t)$ given in Eq. (43)
Eigenvalue equation $\hat{H}_{0}\Phi_{n}^{(0)}(x,t)=\hbar\omega_{0}\left(n+1/2\right)\Phi_{n}^{(0)}(x,t)$ $\hat{I}(t)\Phi_{n}(x,t)=\hbar\left(n+1/2\right)\Phi_{n}(x,t)$
Eigenfunction $\Phi_{n}^{(0)}(x,t)$ given by Eq. (18) $\Phi_{n}(x,t)$ given by Eq. (59)
General solution $\Psi^{(0)}(x,t)=\sum_{n=0}^{\infty}C_{n}^{(0)}\exp[i\alpha_{n}^{(0)}(t)]\Phi_{n}^{(0)}(x,t)$ $\Psi(x,t)=\sum_{n=0}^{\infty}C_{n}\exp[i\alpha_{n}(t)]\Phi_{n}(x,t)$
[…]
Table 2: Comparison between the time-independent and time-dependent quantum harmonic oscillators.
IV.2 The time-dependent quantum harmonic oscillator and the generation of squeezed states
It is well known that the wave functions of quantum harmonic oscillators with time-dependent parameters describe the so-called squeezed states [61, 30, 31, 45, 96, 97]. One has a squeezed state when the variance of an observable in this state becomes smaller than its value for a coherent state, leading to an increase in the variance of the complementary observable in such a way that the Heisenberg uncertainty principle is still satisfied [98, 96]. The parameters that characterize when the oscillator states are squeezed, namely the squeezing parameter $r(t)$ and the squeezing phase $\varphi(t)$, can be defined as (for more details, see Refs. [99, 100, 66])
$r(t)=\cosh^{-1}\{[\lambda(t)]^{\frac{1}{2}}\},$ (76)
where
$\lambda(t)=\frac{m_{0}^{2}\dot{\rho}(t)^{2}+\rho(t)^{-2}+m_{0}^{2}\omega(t)^{2}\rho(t)^{2}+2m_{0}\omega(t)}{4m_{0}\omega(t)},$ (77)
and
$\varphi(t)=\cos^{-1}\left\{\frac{1+m_{0}\omega(t)\rho(t)^{2}-2\cosh^{2}[r(t)]}{2\sinh[r(t)]\cosh[r(t)]}\right\}.$ (78)
Note that, when the oscillator is time-independent, the parameter $\rho(t)$ is given by Eq. (75) and when substituting it into Eqs. (76)-(78), one obtains that the squeezing parameter is null and the squeezing phase is undefined. Thus, for a time-independent harmonic oscillator, there are no squeezed states.
Taking into account the wave function $\Psi_{n}(x,t)$, we can define the expected value of a given observable $\hat{{\cal O}}$ in the state $\Psi_{n}(x,t)$ for any $t\geq 0$ as [7, 9]
[…]
Using this equation and the definition of the variance associated with the operator $\hat{{\cal O}}$, namely $\sigma_{\hat{{\cal O}}}^{2}(n,t)=[\langle\hat{{\cal O}}^{2}\rangle(n,t)]-[\langle\hat{{\cal O}}\rangle(n,t)]^{2}$, we have, for the position and momentum operators, that [66]
$\sigma_{\hat{x}}^{2}(n,t)\;=$$\;\,\bigl{\{}\cosh^{2}[r(t)]+2\sinh[r(t)]\cosh[r(t)]\cos[\varphi(t)]+\sinh^{2}[r(t)]\bigr{\}}\biggl{(}n+\frac{1}{2}\biggr{)}\frac{\hbar}{m_{0}\omega(t)},$ (80)
$\sigma_{\hat{p}}^{2}(n,t)\;=$$\;\,\bigl{\{}\cosh^{2}[r(t)]-2\sinh[r(t)]\cosh[r(t)]\cos[\varphi(t)]+\sinh^{2}[r(t)]\bigr{\}}\biggl{(}n+\frac{1}{2}\biggr{)}\hbar m_{0}\omega(t).$ (81)
As discussed, in the time-independent case [$\omega(t)=\omega_{0}$] we have $r(t)=0$, so that Eqs. (80) and (81) recover the usual variances of a time-independent harmonic oscillator [7].
The system subjected to a TDHO potential can also make transitions between different states. In this sense, the probability ${\cal P}(t)_{0\to\nu}$ of the oscillator making a transition from its ground state to an excited state $\nu$ is defined by [101, 96, 98]
${\cal P}(t)_{0\to\nu}=\frac{\nu!\tanh^{\nu}[r(t)]}{2^{\nu}[(\nu/2)!]^{2}\cosh[r(t)]},$ (82)
where $\nu=0,2,4,...\;$. When we consider $\nu=0$ in this equation, one has the probability of persistence in the fundamental state, ${\cal P}_{p}(t)$. With this, one can also obtain the probability of excitation, which is given by ${\cal P}_{e}(t)=1-{\cal P}_{p}(t)$ [96].
IV.3 Application to an undergraduate and postgraduate problem
In this section, by means of the results presented in previous sections, we solve the problem 2.14 of Ref. [9] (or the problem 11.19 of Ref. [11]), which reads as follows: “A particle is in the ground state of the harmonic oscillator with classical frequency $\omega_{0}$, when suddenly the spring constant quadruples, so $\omega_{1}=2\omega_{0}$, without initially changing the wave function. What is the probability that a measurement of the energy would still return the value $\hbar\omega_{0}/2$? What is the probability of getting $\hbar\omega_{0}$?”
Let us start discussing this problem considering a general value for $\omega_{1}$. In a general way, this problem can be modeled as a TDHO with a time-dependent frequency defined by
$\omega(t)=\begin{cases}\omega_{0},&t<0,\\ \omega_{1},&t\geq 0.\end{cases}$ (83)
In this case, the energy spectrum in the interval $t\geq 0$ simply becomes $E_{n}^{\prime}=\hbar\omega_{1}\left(n+1/2\right)$. Besides this, as discussed, quantum harmonic oscillators with time-dependent parameters describe squeezed states. To calculate the squeezing parameter $r(t)$, we first have to calculate the parameter $\rho(t)$, which is calculated from the Ermakov-Pinney equation [Eq. (42)]. One can find that the solution of the Ermakov-Pinney equation for the interval $t\geq 0$, and subjected to the initial conditions $\rho(t)|_{t=0}=\rho_{0}$ and $\dot{\rho}(t)|_{t=0}=0$, is given by [65]
$\rho(t)=\sqrt{\frac{\omega_{0}\sin^{2}(\omega_{1}t)}{m_{0}\omega_{1}^{2}}+\frac{\cos^{2}(\omega_{1}t)}{m_{0}\omega_{0}}}.$ (84)
Substituting this result into Eqs. (76) and (77), we obtain
$r(t)=\cosh^{-1}\left(\frac{\omega_{0}+\omega_{1}}{2\sqrt{\omega_{0}\omega_{1}}}\right).$ (85)
Moreover, since the system is initially in the ground state, it is also interesting to calculate the probability of its persistence in the ground state, which is obtained from Eq. (82) with $\nu=0$ and is written as
${\cal P}_{p}(t)=\frac{1}{\cosh[r(t)]}.$ (86)
Substituting Eq. (85) in this equation, one obtains
${\cal P}_{p}(t)=\frac{2\sqrt{\omega_{0}\omega_{1}}}{\omega_{0}+\omega_{1}}.$ (87)
Note that in both limits $\omega_{1}/\omega_{0}\to 0$ and $\omega_{1}/\omega_{0}\to\infty$, we have $P_{p}(t)\to 0$, which means that after the system changes its frequency in these limit situations, it will also be excited. The behavior of ${\cal P}_{p}(t)$ given by Eq. (87) is illustrated in Fig. 1.
Figure 1: Behavior of $P_{p}(t)$ given by Eq. (87), as a function of $\omega_{1}/\omega_{0}$.
For the problem 2.14 of Ref. [9] (or the problem 11.19 of Ref. [11]), we have $\omega_{1}=2\omega_{0}$. In this case, the energy spectrum in the interval $t\geq 0$ becomes $E_{n}^{\prime}=2\hbar\omega_{0}\left(n+1/2\right)$. Note that, in this interval, the energy associated to the ground state $(n=0)$ is $E_{0}^{\prime}=\hbar\omega_{0}$. This means that, when $t\geq 0$, the probability of a measurement of the energy resulting in any value less than $E_{0}^{\prime}$ is zero and, thus, such probability for $\hbar\omega_{0}/2$ is zero. On the other hand, the probability of getting $\hbar\omega_{0}$ may not be zero and is simply given by Eq. (87) with $\omega_{1}=2\omega_{0}$ (since $\hbar\omega_{0}$ is the energy corresponding to the ground state for $t\geq 0$) which results in
${\cal P}_{p}(t)=\frac{2\sqrt{2}}{3}\cong 0.943.$ (88)
This result is the same as that obtained from the elementary quantum mechanics methods discussed in undergraduate and postgraduate courses (see Refs. [9, 11]). Next, we analyze a model where the usual methods taught at undergraduate and postgraduate level cannot be applied directly.
IV.4 Quantum dynamics of a particle in a Paul trap
A Paul trap is a device used to capture and manipulate charged particles, such as ions, using oscillating electromagnetic fields, and has applications in different areas of physics, such as cooling processes and mass spectroscopy [102, 37]. In this sense, due to the importance of this device for the modern physics, it could be very enriching to make a basic discussion on this topic in undergraduate and postgraduate quantum mechanics courses. However, the tools usually taught in these courses are not enough to deal with this problem, since the dynamics of a particle in a Paul trap is effectively modeled by a TDHO with time-dependent frequency of the form [37, 92]
$\omega(t)=\frac{\omega_{0}}{\sqrt{\beta+\gamma}}\sqrt{\beta+\gamma\cos\left(\frac{2\pi t}{\tau}\right)},$ (89)
where $\tau$ is the oscillation period [thus, $\omega(t+\tau)=\omega(t)$], and $\beta$ and $\gamma$ are numerical factors such that $\beta>\gamma$. Using the LR method, this problem can be solved exactly and, thus, it can be interesting to make a brief analyzes of this case (something similar has already been done, for example, in Ref. [92]). The solution of the Ermakov-Pinney equation for $t\geq 0$, with initial conditions $\rho(t)|_{t=0}=\rho_{0}$ and $\dot{\rho}(t)|_{t=0}=0$, is given by
$\rho(t)=\sqrt{\frac{M_{C}\left[\frac{\beta\omega_{0}^{2}\tau^{2}}{\pi^{2}(\beta+\gamma)},-\frac{\gamma\omega_{0}^{2}\tau^{2}}{2\pi^{2}(\beta+\gamma)},\frac{\pi t}{\tau}\right]^{2}}{m_{0}\omega_{0}M_{C}\left[\frac{\beta\omega_{0}^{2}\tau^{2}}{\pi^{2}(\beta+\gamma)},-\frac{\gamma\omega_{0}^{2}\tau^{2}}{2\pi^{2}(\beta+\gamma)},0\right]^{2}}+\frac{\omega_{0}\tau^{2}M_{S}\left[\frac{\beta\omega_{0}^{2}\tau^{2}}{\pi^{2}(\beta+\gamma)},-\frac{\gamma\omega_{0}^{2}\tau^{2}}{2\pi^{2}(\beta+\gamma)},\frac{\pi t}{\tau}\right]^{2}}{\pi^{2}m_{0}\dot{M}_{S}\left[\frac{\beta\omega_{0}^{2}\tau^{2}}{\pi^{2}(\beta+\gamma)},-\frac{\gamma\omega_{0}^{2}\tau^{2}}{2\pi^{2}(\beta+\gamma)},\frac{\pi t}{\tau}\right]^{2}\Bigr{|}_{t=0}}},$ (90)
in which $M_{C}$ and $M_{S}$ are the even and odd Mathieu functions, respectively. Using Eqs. (89) and (90) in Eqs. (76) and (77), we obtain the squeezing parameter $r(t)$ for this model, whose behavior is shown in Fig. 2.
Figure 2: Behavior of the squeezing parameter $r(t)$ associated with the model defined by Eq. (89), as a function of $t/\tau$. In this figure we consider, in arbitrary units, $\beta=1$, $\gamma=1/2$ and $\omega_{0}\tau=3$.
Note the quasi-periodic oscillatory behavior of the squeezing parameter $r(t)$ [92], which occurs naturally due to the form of the frequency $\omega(t)$ in Eq. (89). This also influences the behavior of all the other associated quantities [see Eqs. (80)-(82)], as can be seen in Fig. 3 for the variances of position and momentum, and in Fig. 4 for the probability of persistence in the ground state ${\cal P}_{p}(t)$ and the probability of excitation ${\cal P}_{e}(t)$.
Figure 3: Behavior of the variances (a) $\sigma_{\hat{x}}^{2}(n,t)/\overline{\sigma}_{\hat{x}}^{2}(n)$ and (b) $\sigma_{\hat{p}}^{2}(n,t)/\overline{\sigma}_{\hat{p}}^{2}(n)$, with $\overline{\sigma}_{\hat{x}}^{2}(n)=\tfrac{\hbar(n+{1}/{2})}{m_{0}\omega_{0}}$ and $\overline{\sigma}_{\hat{p}}^{2}(n)=\hbar m_{0}\omega_{0}(n+{1}/{2})$, as a function of $t/\tau$. In these figures we consider, in arbitrary units, $\beta=1$, $\gamma=1/2$ and $\omega_{0}\tau=3$.
Figure 4: Behavior of the probabilities (a) ${\cal P}_{p}(t)$ and (b) ${\cal P}_{e}(t)$, as a function of $t/\tau$. In these figures we consider, in arbitrary units, $\beta=1$, $\gamma=1/2$ and $\omega_{0}\tau=3$.
V Final Remarks
In Sec. II, we briefly reviewed the algebraic method to obtain the solution of the Schrödinger equation for the problem of a time-independent quantum harmonic oscillator, which is widely studied in quantum mechanics courses [9, 7, 8]. This section prepares the reader for the subsequent sections, since the application of the Lewis-Riesenfeld method to a time-dependent harmonic oscillator prescribes a sequence of steps similar to those already used in the time-independent case (and which were summarized in the end of Sec. II).
In Sec. III, we discussed the general formulation of the Lewis-Riesenfeld method [12, 13, 16], step-by-step and with a didactic approach, and we presented a summary of the method, organized into four steps. We highlighted that the usual way to approach the problem of the time-independent quantum harmonic oscillator (looking for eigenfunctions of the time-independent Hamiltonian operator) is a particular case of the Lewis-Riesenfeld approach (looking for eigenfunctions of the Lewis’ invariant operator). We also presented Table 1, where we summarized a comparison between the characteristics of a general approach to obtain the solution of a time-independent case and the correspondent ones of a time-dependent situation.
In Sec. IV.1, we provided a didactic application of the Lewis-Riesenfeld method, obtaining the wave function for the case of a time-dependent harmonic oscillator with a generic time-dependent frequency. We highlighted that the steps for finding the eigenstates and eigenvalues of the Lewis’ invariant operator for this system (Lewis’ invariant operator is written in terms of ladder operators) are similar to those of the usual algebraic method used to find the eigenstates and eigenvalues of the Hamiltonian of a time-independent quantum harmonic oscillator (the time-invariant Hamiltonian operator is written in terms of ladder operators). Closing this section, we presented Table 2, where we summarized a comparison between the characteristics of the time-independent oscillator addressed via algebraic approach (discussed in Sec. II) and the correspondent ones of a time-dependent oscillator addressed via Lewis-Riesenfeld method.
In Sec. IV.2, we briefly discussed the relation between the quantum time-dependent harmonic oscillator and the generation of squeezed states, using, in Sec. IV.3, these concepts to calculate the transition probability for a time-dependent harmonic oscillator experiencing a sudden jump in its frequency [9, 11]. In Sec. IV.4, we also used all these tools to briefly study the dynamics of a quantum particle in a Paul trap [37, 92]. These calculations performed in Secs. IV.3 and IV.4 can be used as didactic examples of applications of the Lewis-Riesenfeld method, which exactly solves the Schrödinger equation with explicitly time-dependent Hamiltonians, in quantum mechanics courses.
The presented discussion aims to contribute to the effort already begun in articles in physics teaching journals [10, 15, 33, 67, 68] to improve the discussions on time-dependent quantum harmonic oscillators in the context of undergraduate courses.
Declaration of competing interest
The authors declare no conflict of interest.
Acknowledgements
S.S.C. and L.Q. were supported by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES), Finance Code 001. This work was partially supported by CNPq - Brazil, Processo 408735/2023-6 CNPq/MCTI.
Appendix A Temporal independence of the eigenvalues $\lambda$
Applying the operator $\partial\hat{I}(t)/\partial t$ in the state $|\lambda,\zeta;t\rangle$,
$\frac{\partial\hat{I}(t)}{\partial t}|\lambda,\zeta;t\rangle=-\frac{1}{i\hbar}\bigl{[}\hat{I}(t)\hat{H}(t)-\hat{H}(t)\hat{I}(t)\bigr{]}|\lambda,\zeta;t\rangle,$ (91)
acting $\langle\lambda^{\prime},\zeta^{\prime};t|$ in Eq. (91) and using Eq. (24), we obtain
$\langle\lambda^{\prime},\zeta^{\prime};t|\frac{\partial\hat{I}(t)}{\partial t}|\lambda,\zeta;t\rangle=-\frac{\left(\lambda^{\prime}-\lambda\right)}{i\hbar}\langle\lambda^{\prime},\zeta^{\prime};t|\hat{H}(t)|\lambda,\zeta;t\rangle.$ (92)
However, deriving Eq. (24) with respect to time,
$\frac{\partial\hat{I}(t)}{\partial t}\bigl{|}\lambda,\zeta;t\rangle+\hat{I}(t)\frac{\partial}{\partial t}\bigl{|}\lambda,\zeta;t\rangle=\frac{\partial\lambda}{\partial t}\bigl{|}\lambda,\zeta;t\rangle+\lambda\frac{\partial}{\partial t}\bigl{|}\lambda,\zeta;t\rangle,$ (93)
operating $\langle\lambda^{\prime},\zeta^{\prime};t|$ in Eq. (93), utilizing the Eq. (26) and noting that the operator $\hat{I}(t)$ is Hermitian, we find
$\langle\lambda^{\prime},\zeta^{\prime};t|\frac{\partial\hat{I}(t)}{\partial t}|\lambda,\zeta;t\rangle=\frac{\partial\lambda}{\partial t}\delta_{\lambda^{\prime},\lambda}\delta_{\zeta^{\prime},\zeta}-\left(\lambda^{\prime}-\lambda\right)\langle\lambda^{\prime},\zeta^{\prime};t|\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle.$ (94)
Since Eqs. (92) and (94) are equal, consequently
$\frac{\partial\lambda}{\partial t}\delta_{\lambda^{\prime},\lambda}\delta_{\zeta^{\prime},\zeta}-\left(\lambda^{\prime}-\lambda\right)\langle\lambda^{\prime},\zeta^{\prime};t|\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle+\frac{\left(\lambda^{\prime}-\lambda\right)}{i\hbar}\langle\lambda^{\prime},\zeta^{\prime};t|\hat{H}(t)|\lambda,\zeta;t\rangle=0.$ (95)
Hence, for $\lambda^{\prime}=\lambda$ and $\zeta^{\prime}=\zeta$, it will result from Eq. (95) that
$\frac{\partial\lambda}{\partial t}=0.$ (96)
Appendix B The relation between the eigenstates of $\hat{I}(t)$ and $\hat{H}(t)$ operators
According to Eq. (23), the action of the operator $\hat{I}(t)$ on any state ket, which is a solution of the Schrödinger equation, produces a new solution of it. From Eqs. (93) and (96), we have
$\frac{\partial\hat{I}(t)}{\partial t}|\lambda,\zeta;t\rangle=\bigl{[}\lambda-\hat{I}(t)\bigr{]}\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle,$ (97)
then, multiplying Eq. (97) by bra $\langle\lambda^{\prime},\zeta^{\prime};t|$, we get
$\langle\lambda^{\prime},\zeta^{\prime};t|\frac{\partial\hat{I}(t)}{\partial t}|\lambda,\zeta;t\rangle=\langle\lambda^{\prime},\zeta^{\prime};t|\bigl{[}\lambda-\hat{I}(t)\bigr{]}\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle,$ (98)
and, equating Eqs. (92) and (98), we obtain
$i\hbar\left(\lambda-\lambda^{\prime}\right)\langle\lambda^{\prime},\zeta^{\prime};t|\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle=\left(\lambda-\lambda^{\prime}\right)\langle\lambda^{\prime},\zeta^{\prime};t|\hat{H}(t)|\lambda,\zeta;t\rangle.$ (99)
In the general case where $\lambda^{\prime}\neq\lambda$ and $\zeta^{\prime}\neq\zeta$, we find
$i\hbar\langle\lambda^{\prime},\zeta^{\prime};t|\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle=\langle\lambda^{\prime},\zeta^{\prime};t|\hat{H}(t)|\lambda,\zeta;t\rangle.$ (100)
To analyze the case $\lambda^{\prime}=\lambda$ and $\zeta^{\prime}=\zeta$, we will proceed as follows. Knowing that multiplying the state vector by a phase factor of the form $\exp[i\alpha_{\lambda,\zeta}(t)]$ [$\alpha_{\lambda,\zeta}(t)$ is a real and time-dependent function] will not change the probability amplitude because the scalar product will remain unchanged, so we can define a new set of state vectors, which relate to the initial state ket $|\lambda,\zeta;t\rangle$ by means of the gauge transformation [16]
$|\lambda,\zeta;t\rangle_{\alpha}=\exp[i\alpha_{\lambda,\zeta}(t)]|\lambda,\zeta;t\rangle.$ (101)
According to Ref. [16], because the operator $\hat{I}(t)$ contains no temporal derivatives, one can immediately conclude that
$\hat{I}(t)|\lambda,\zeta;t\rangle_{\alpha}=\lambda|\lambda,\zeta;t\rangle_{\alpha},$ (102)
that is, the states $|\lambda,\zeta;t\rangle_{\alpha}$ are also eigenstates of the operator $\hat{I}(t)$, so that $|\lambda,\zeta;t\rangle_{\alpha}$ and $|\lambda,\zeta;t\rangle$ share the same eigenvalue spectrum and, consequently, the same completeness and orthonormality properties. Thus, if we require that the eigenstates $|\lambda,\zeta;t\rangle_{\alpha}$ satisfy the Schrödinger equation:
$i\hbar\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle_{\alpha}=\hat{H}(t)|\lambda,\zeta;t\rangle_{\alpha},$ (103)
then it will be possible to find the phases $\alpha_{\lambda,\zeta}(t)$. For this, deriving Eq. (101) with respect to time and substituting the result in Eq. (103), we arrive at
$i\frac{\partial}{\partial t}|\lambda,\zeta;t\rangle-\frac{d\alpha_{\lambda,\zeta}(t)}{dt}|\lambda,\zeta;t\rangle=\frac{1}{\hbar}\hat{H}(t)|\lambda,\zeta;t\rangle.$ (104)
Next, multiplying Eq. (104) by bra $\langle\lambda^{\prime},\zeta^{\prime};t|$ and using Eq. (26), we have
$\frac{d\alpha_{\lambda,\zeta}(t)}{dt}\delta_{\lambda^{\prime},\lambda}\delta_{\zeta^{\prime},\zeta}=\langle\lambda^{\prime},\zeta^{\prime};t|\biggl{[}i\frac{\partial}{\partial t}-\frac{1}{\hbar}\hat{H}(t)\biggr{]}|\lambda,\zeta;t\rangle.$ (105)
When $\lambda^{\prime}\neq\lambda$ and $\zeta^{\prime}\neq\zeta$, Eq. (105) will be [see Eq. (100)]
$\langle\lambda^{\prime},\zeta^{\prime};t|\left[i\frac{\partial}{\partial t}-\frac{1}{\hbar}\hat{H}(t)\right]|\lambda,\zeta;t\rangle=0,$ (106)
while for $\lambda^{\prime}=\lambda$ and $\zeta^{\prime}=\zeta$, Eq. (105) will fall to
$\frac{d\alpha_{\lambda,\zeta}(t)}{dt}=\langle\lambda,\zeta;t|\left[i\frac{\partial}{\partial t}-\frac{1}{\hbar}\hat{H}(t)\right]|\lambda,\zeta;t\rangle,$ (107)
whose formal solution is given by
[…]
Having the phase functions in hand, we can determine the eigenstates $|\lambda,\zeta;t\rangle_{\alpha}$ since, in principle, we know how to calculate the eigenstates $|\lambda,\zeta;t\rangle$ from Eq. (24). Therefore, with the eigenstates $|\lambda,\zeta;t\rangle_{\alpha}$ being solutions of the Schrödinger equation, any other solution can be described as a linear combination of these states. Accordingly, given that $|\psi(t)\rangle$ is a solution of the Schrödinger equation, then we can describe $|\psi(t)\rangle$ as a superposition of $|\lambda,\zeta,t\rangle_{\alpha}$ in the form
$|\psi(t)\rangle=\sum_{\lambda,\zeta}C_{\lambda,\zeta}\exp\left[i\alpha_{\lambda,\zeta}(t)\right]|\lambda,\zeta;t\rangle,$ (109)
where the coefficients $C_{\lambda,\zeta}$ depend on the initial conditions and are given by
$C_{\lambda,\zeta}=\langle\lambda,\zeta;0|\psi\left(0\right)\rangle.$ (110)
Appendix C Main steps to obtain Eq. (59)
Introducing the unitary operator $\hat{{\cal U}}(t)$, defined by [30, 31]
$\hat{{\cal U}}(t)=\exp\left[-\frac{im(t)\dot{\rho}(t)\hat{x}^{2}}{2\hbar\rho(t)}\right],$ (111)
we can rewrite Eq. (57), in position space, as
$\Phi_{0}(x,t)=\hat{{\cal U}}^{\dagger}(t)\biggl{[}\frac{1}{\pi\hbar\rho(t)^{2}}\biggr{]}^{\frac{1}{4}}\exp\left[-\frac{x^{2}}{2\hbar\rho(t)^{2}}\right].$ (112)
Defining
$\Phi_{0}^{\prime}(x,t)=\biggl{[}\frac{1}{\pi\hbar\rho(t)^{2}}\biggr{]}^{\frac{1}{4}}\exp\left[-\frac{x^{2}}{2\hbar\rho(t)^{2}}\right],$ (113)
and substituting the above equation into Eq. (112), we have
$\Phi_{0}(x,t)=\hat{{\cal U}}^{\dagger}(t)\Phi_{0}^{\prime}(x,t).$ (114)
Furthermore, using the Baker-Campbell-Hausdorff relation [7, 96]:
$\exp(\hat{{\cal G}})\hat{{\cal O}}\exp(-\hat{{\cal G}})=\hat{{\cal O}}+[\hat{{\cal G}},\hat{{\cal O}}]+\frac{1}{2!}[\hat{{\cal G}},[\hat{{\cal G}},\hat{{\cal O}}]]+\ldots+\frac{1}{k!}[\hat{{\cal G}},[\hat{{\cal G}},[\hat{{\cal G}},\ldots[\hat{{\cal G}},\hat{{\cal O}}]]]\ldots]+\ldots,$ (115)
we have
$\hat{a}^{\dagger}(t)=\hat{{\cal U}}^{\dagger}(t)\hat{a}^{\prime\dagger}(t)\,\hat{{\cal U}}(t),$ (116)
where (remember that we are in position space)
$\hat{a}^{\prime\dagger}(t)=\frac{1}{\sqrt{2\hbar}}\left[\frac{x}{\rho(t)}-\hbar\rho(t)\frac{\partial}{\partial x}\right].$ (117)
As such, by analogy with the time-independent harmonic oscillator, we have
$\Phi_{n}^{\prime}(x,t)=\frac{\bigl{[}\hat{a}^{\prime\dagger}(t)\bigr{]}^{n}}{\sqrt{n!}}\,\Phi_{0}^{\prime}(x,t).$ (118)
Consequently, using Eqs. (113) and (117) in Eq. (118), we find
$\Phi_{n}^{\prime}(x,t)=\frac{1}{\sqrt{2^{n}n!}}\biggl{[}\frac{1}{\pi\hbar\rho(t)^{2}}\biggr{]}^{\frac{1}{4}}\left[\frac{x}{\hbar^{\frac{1}{2}}\rho(t)}-\hbar^{\frac{1}{2}}\rho(t)\frac{\partial}{\partial x}\right]^{n}\exp\left[-\frac{x^{2}}{2\hbar\rho(t)^{2}}\right].$ (119)
However, knowing the following relation involving the Hermite polynomials [94]:
${\cal H}_{n}(z)=\exp\left(\frac{z^{2}}{2}\right)\left(z-\frac{\partial}{\partial z}\right)^{n}\exp\left(-\frac{z^{2}}{2}\right),$ (120)
so, by changing variables $z=x/[\hbar^{\frac{1}{2}}\rho(t)]$, we get
${\cal H}_{n}\left[\frac{x}{\hbar^{\frac{1}{2}}\rho(t)}\right]=\exp\left[\frac{x^{2}}{2\hbar\rho(t)^{2}}\right]\left[\frac{x}{\hbar^{\frac{1}{2}}\rho(t)}-\hbar^{\frac{1}{2}}\rho(t)\frac{\partial}{\partial x}\right]^{n}\exp\left[-\frac{x^{2}}{2\hbar\rho(t)^{2}}\right].$ (121)
Hence, rewriting Eq. (119) as
$\Phi_{n}^{\prime}(x,t)=\frac{1}{\sqrt{2^{n}n!}}\biggl{[}\frac{1}{\pi\hbar\rho(t)^{2}}\biggr{]}^{\frac{1}{4}}\left\{\exp\left[\frac{x^{2}}{2\hbar\rho(t)^{2}}\right]\left[\frac{x}{\hbar^{\frac{1}{2}}\rho(t)}-\hbar^{\frac{1}{2}}\rho(t)\frac{\partial}{\partial x}\right]^{n}\exp\left[-\frac{x^{2}}{2\hbar\rho(t)^{2}}\right]\right\}\exp\left[-\frac{x^{2}}{2\hbar\rho(t)^{2}}\right],$ (122)
substituting Eq. (121) into Eq. (122) and returning to the original function by [30, 31]
$\Phi_{n}(x,t)=\hat{{\cal U}}^{\dagger}(t)\Phi_{n}^{\prime}(x,t),$ (123)
finally
$\Phi_{n}(x,t)=\frac{1}{\sqrt{2^{n}n!}}\Phi_{0}(x,t){\cal H}_{n}\left[\frac{x}{\hbar^{\frac{1}{2}}\rho(t)}\right].$ (124)
Appendix D Main steps to obtain Eq. (64)
Noting that we can write the scalar product between two state vectors, $|A\rangle$ and $|B\rangle$, in position space as [7]
[…]
making
$|A\rangle=\frac{\partial}{\partial t}|0;t\rangle,\;\;\langle B|=\langle 0;t|,$ (126)
we arrive in
[…]
However, realizing that
$\langle x|\frac{\partial}{\partial t}|0;t\rangle=\frac{\partial\Phi_{0}(x,t)}{\partial t},\;\;\langle 0;t|x\rangle=\Phi_{0}^{*}(x,t),$ (128)
the Eq. (127) will become
[…]
Given this, rewriting Eq. (57) as
$\Phi_{0}(x,t)=b(t)\exp\left[-c(t)x^{2}\right],$ (130)
where
$b(t)=\biggl{[}\frac{1}{\pi\hbar\rho(t)^{2}}\biggr{]}^{\frac{1}{4}},\;\;c(t)=-\frac{im_{0}}{2\hbar}\biggl{[}\frac{\dot{\rho}(t)}{\rho(t)}+\frac{i}{m_{0}\rho(t)^{2}}\biggr{]},$ (131)
deriving Eq. (130) with respect to time, we find
$\frac{\partial\Phi_{0}(x,t)}{\partial t}=\left[F(t)x^{2}-\frac{\dot{\rho}(t)}{2\rho(t)}\right]b(t)\exp\left[-c(t)x^{2}\right],$ (132)
in which
$F(t)=\frac{im_{0}\ddot{\rho}(t)}{2\hbar\rho(t)}-\frac{im_{0}\dot{\rho}(t)^{2}}{2\hbar\rho(t)^{2}}+\frac{\dot{\rho}(t)}{\hbar\rho(t)^{3}},$ (133)
so, substituting the complex conjugate of Eq. (130) and Eq. (132) into Eq. (129), we arrive at
[…]
Solving the Gaussian integrals and returning to the original variables, finally [16]
$\langle 0;t|\frac{\partial}{\partial t}|0;t\rangle=\frac{im_{0}}{4}[\rho(t)\ddot{\rho}(t)-\dot{\rho}(t)^{2}].$ (135)